In the equation V^n=K.a^j.x ,what number must n & j be to make the equation dimensionally correct? K= a dimensionless number, x = L , v= velocity and a= acceleration.
1
Expert's answer
2019-03-14T17:37:13-0400

The given equation is:

"v^n = ka^jx."

Let's start by writing the known dimensions (units) of the values:

"[k] = 1;~[x] = m;~[v] = \frac{m}{s};~[a] = \frac{m}{s^2}."

Now let's rewrite the equation using only the units (e.g. assuming each value is equal to unit value) and rework the powers:

"(\frac{m}{s})^n = (\frac{m}{s^2})^jm; \\
m^ns^{-n} = m^{j+1}s^{-2j}."

Now, for the equation to be correct, the powers of each unit (i.e. meters and seconds) have to match correspondingly:

"\begin{cases}
 n = j + 1; \\
 -n = -2j.
\end{cases}"

Now it's easy to solve these equations together for n and j:

"2j = j + 1; \implies j = 1; \\ 
n = 1 + 1 = 2."

Therefore the correct form of the initial equation, with the found n and j is:

"v^2 = kax."